Dawson's Repair Service orders parts from an electronic company, which advertises its parts to be no more than 11% defective. What is the probability that Bill Dawson finds 12 or more parts out of a sample of 100 to be defective?
Answer by Boreal(15235)

Find the probability that 11, 10, 9,... are defective.
This can be done using the binomial distribution.
for 11 it is 100C11*.11^11*.89^89=0.1265
for 10 it is 100C10*.11^10*.89^90=0.1251
for 9 it is 0.1112
for 8 it is 0.0881
for 7, 0.0613
etc. down to 0
OR, can do binomcdf(100,.11,11) which will get everything from the left side to 11. That probability is 0.5794.
The answer is the complement or 1-0.5794=0.4206. (Also 1-binomcdf(100,0.11,11))
-
can check with normal approximation
np=11
np(1-p)=V=11(.89)=9.79
sqrt(V)=sd=3.13
z > (11.5-11)/3.13, with continuity correction factor >0.5/3.13 or 0.16. That probability is 0.4364, which is a reasonable approximation of the actual answer.